OpenStudy (anonymous):

Complete the equation that is related to the given equation. x – (–6.9) = 42.8 x = 42.8 +

OpenStudy (anonymous):

Complete the equation that is related to the given equation. x – (–6.9) = 42.8 x = 42.8 +

sammixboo (sammixboo):

I think it is x = 42.8 + (-6.9) Because if you are going x - (-6.9) = 42.8 your first step is to add -6.9 to both sides

sammixboo (sammixboo):

Understand @funjjademelb65